OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/6

OpenStudy (mathmate):

Can you prove the case when n=1?

OpenStudy (anonymous):

@mathmate no, i honestly do not understand how to do :(

OpenStudy (mathmate):

The process of mathematical induction is a three step process: 1. prove a base case, usually when n=0 or n=1. 2. ASSUMING that the conjecture is true for the general case n, prove that it is true for the case n+1. 3. If 1 and 2 are both true, then the conjecture is true for all n where \( n\in N \).

OpenStudy (anonymous):

@mathmate do you use the equation right of the equal sign to figure out if it is true?

OpenStudy (mathmate):

The given conjecture is 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/6 should be true for all positive integers. So let's do the first step of trying n=1. Can you do that?

OpenStudy (mathmate):

You will need to evaluate the left-hand side and right-hand side according to the equation. The left-hand side has n terms, while the right hand side has only one term.

OpenStudy (mathmate):

For n=1, both sides have only one term.

OpenStudy (anonymous):

@mathmate oh so the left and right equations have to equal each other?

OpenStudy (mathmate):

If the conjecture is true, they should be equal. If they are not equal for ANY value of n, the conjecture is not true.

OpenStudy (mathmate):

Can you start by evaluating the left-hand side for n=1?

OpenStudy (anonymous):

@mathmate oh okay! that makes a lot more sense! so right now when i plug in 1 on each side , i have \[[4(1)][4(1)+2]=\frac{ 4[4(1)+1][8(1)+7] }{ 6 }\] the left side equals \[4\times6\] or 24 and the right side equals \[(4\times5\times15)\div6\] \[300\div6\] which equals 50 so they do not equal each other

OpenStudy (mathmate):

And the conclusion is...... But before we go further, I would always check if the question was posted or understood correctly. Teachers don't usually give a freebie like that.

OpenStudy (anonymous):

@mathmate well since it is false, i have to show why and how

OpenStudy (mathmate):

You already showed a counter example that the conjecture is false for n=1. So the conjecture is false. (Note that, on the contrary, showing that it works for n=1 does not prove that it works for all n). However, as I mentioned, make sure there are no typos.

OpenStudy (mathmate):

... no typos in the question that you posted.

OpenStudy (anonymous):

@mathmate i double checked and everything is correct. should i use another number for show that it is false?

OpenStudy (mathmate):

In mathematics, if you showed one single case (ex. n=1) is false, the conjecture is false. Compare with murder cases before a judge. Often when the prosecutor proves that the witness has lied about a detail in the testimony, the testimony is thrown out almost all the time.

OpenStudy (anonymous):

@mathmate thank you so much for your help , honestly. i finally understand this!! thank you again :) have a great new year!!

OpenStudy (mathmate):

yw! :)